OpenStudy (anonymous):

differentiate y=9e^x(x^.5)

OpenStudy (anonymous):

can you use Equation to enter it ?

zepdrix (zepdrix):

\[\Large\rm y=9x^{1/2}e^x\]Need help differentiating this bunch? :o

OpenStudy (anonymous):

Oh I see. It is much easier now

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Woops D: Worm, you forgot your product rule!

zepdrix (zepdrix):

And I don't know why you have an integral sign in your picture.

OpenStudy (anonymous):

Are we finding y' ?

zepdrix (zepdrix):

She asked the question before ^^ So I think she already got the answer But anyway,\[\Large\rm \left(9x^{1/2}e^x\right)'=\color{royalblue}{\left(9x^{1/2}\right)'}e^x+9x^{1/2}\color{royalblue}{\left(e^x\right)'}\] \[\Large\rm \left(9x^{1/2}e^x\right)'=\color{orangered}{\left(\frac{9}{2}x^{-1/2}\right)}e^x+9x^{1/2}\color{orangered}{\left(e^x\right)}\]Yes, y'. Something like that, yah? :o

OpenStudy (anonymous):

I was heading toward the opposite direction. :(

zepdrix (zepdrix):

<:o